### v2

# Evidential Transformers for Improved Image Retrieval

Abstract
We introduce the Evidential Transformer, an uncertainty-driven transformer model for improved and robust image retrieval. In this paper, we make several contributions to content-based image retrieval (CBIR). We incorporate probabilistic methods into image retrieval, achieving robust and reliable results, with evidential classification surpassing traditional training based on multiclass classification as a baseline for deep metric learning. Furthermore, we improve the state-of-the-art retrieval results on several datasets by leveraging the Global Context Vision Transformer (GC ViT) architecture. Our experimental results consistently demonstrate the reliability of our approach, setting a new benchmark in CBIR in all test settings on the Stanford Online Products (SOP) and CUB-200-2011 datasets.
1 Introduction
Content-based image retrieval is a well-known computer vision problem [37, 9]. This problem aims to retrieve images from a database that are visually similar to a given query image. The similarity criterion used here is often measured between the vector representations of images. Such representations are generally sparse and relatively low-dimensional. They capture the semantic content of images by obtaining vector embeddings that contain local and global context information of images, and are vital to building an efficient image retrieval system [1, 6].
Classical well-known pipelines for image retrieval tasks such as [2, 29] begin by representing images using widely-adopted SIFT descriptors [25] as image representations. Similarity between the query and database images is then assessed using metrics like cosine similarity, enabling the retrieval of visually similar images. Nevertheless, with the massive success of convolutional neural networks (CNNs) [19], deep convolutional features have largely supplanted SIFT features due to their state-of-the-art results on several computer vision problems [34, 7, 3, 30, 21, 20, 22, 18]. Following these developments, CNN-based approaches such as [3, 4, 28] leverage neural codes for image retrieval [^1] A neural code is a vector representation of an image, produced by the top layer of a CNN.. Such neural codes perform competitively even when the code-producing network is trained on a task not directly related to image retrieval, like image classification [42]. Prominent methods in this direction, such as [31, 32], fine-tune CNNs for image retrieval without human annotation.
Recently, Vision Transformer (ViT) architectures have outperformed CNNs across a variety of computer vision tasks [11, 23]. In particular, methods using the ViT’s outputs corresponding to the CLS-token as image descriptors for image retrieval have shown state-of-the-art results [12], demonstrating that the ViT is capable of producing more informative embeddings, yielding superior results on multiple benchmark datasets, including CUB-200-2011 [41], SOP [38], and InShop [24]. Reranking methods that utilize attention have also emerged [39]. A recent extension, GC ViT [16], proposed a Vision Transformer architecture that efficiently models local and global contextual relations, producing a feature space where these relations induce a better separability between embeddings of different objects.
However, current methods employ a generic similarity metric to rank retrieved images, which limits their ability to provide high-level retrieval information — specifically, the likelihood that retrieved images closely resemble the query image. Consequently, we propose an evidential learning-driven approach to the transformer model for image retrieval. The idea is that the image similarity must not be confined to image class features but include other features such as the object’s proximity to the camera, local and global scene context, number of pixel objects occupied in the query and database image, etc. These aspects are challenging to model but significantly impact image retrieval effectiveness, and thus, a probabilistic modeling approach can help account for this information in the GC ViT model.
Although there exist many approaches to probabilistic neural networks [15, 14, 5, 13], prior networks and the evidential learning paradigm reduce computational cost relative to ensemble approaches [26]. Therefore, we employ an evidential prior transformer instead of ensemble methods seen in [8, 35, 13], leveraging evidential deep learning [36, 17, 26, 27]. Motivated by recent works demonstrating that classification is a strong baseline for deep metric learning [42], we study the potential of deep evidential classification in providing a more informed baseline for deep metric learning. By setting an evidential prior over the predicted softmax probabilities, we infer image embedding in probabilistic terms.
Subsequently, in this paper, we introduce an approach that efficiently combines attention-based feature maps with uncertainty quantification methods to improve the image-retrieval performance and robustness. Concretely, how embedding uncertainties can be included in the image retrieval process is presented. We observed that uncertainty-driven transformer-based neural representations outperform existing methods in content-based category-level image retrieval. To summarize, our key contributions are:
  • We show that deep evidential classification is a strong baseline for deep metric learning by achieving state-of-the-art results on the CUB-200-2011 dataset.
  • We present a novel, task-agnostic re-ranking method based on uncertainty estimates, which outperforms baseline methods that do not incorporate re-ranking.
  • Our work shows that Dirichlet probability distribution parameters are good neural codes for image retrieval. Additionally, this work introduce a novel continuous embedding method, where each image is mapped into a distribution space and subsequently compared using the Bhattacharyya distance.
  • Our rigorous experiments demonstrate that the our uncertainty-driven tranformer approach significantly outperforms current state-of-the-art image retrieval methods on the CUB-200-2011 and SOP datasets.
2 Methodology
Evidential learning is an effective method for quantifying uncertainty in image retrieval tasks. In contrast to traditional deterministic neural networks that output point predictions, evidential learning explicitly models uncertainty, providing a more robust and informative framework for applications in image retrieval.
The main advantage of evidential networks is that they do not produce a soft-max categorical probability distribution over the class labels but rather a second-order distribution over the soft-max probabilities, allowing them to reason about the uncertainty of the second-order distribution. Such uncertainty can arise from out-of-distribution samples in the dataset, or from images that have embeddings similar to the query but belong to a different class. Assigning a lower rank to these examples could improve the quality of retrieval. This is particularly useful in scenarios where it is crucial to assess the confidence of the model’s predictions, such as in the retrieval of images from large and diverse datasets.
In evidential learning, the output of a neural network is interpreted as a $K$-dimensional parameter vector $\boldsymbol{\alpha}=\left[\alpha_{1},\dots,\alpha_{K}\right]$ of a Dirichlet distribution, which provides a second-order probability distribution over softmax class probabilities $\mathbf{p}$:
[…]
where $K$ is the number of classes in a classification task, and $\mathcal{S}_{K}$ is an K1K-1 simplex:
$\mathcal{S}_{K}=\left\{\mathbf{p}\mid\sum_{i=1}^{K}p_{i}=1\text{ and }0\leq p_{1},\ldots,p_{K}\leq 1\right\}$ (2)
This is a significant departure from the softmax prediction, which only offers a single set of class probabilities $\mathbf{p}$ without any indication of uncertainty. The parameters $\alpha_{k}$ are derived from the evidence $e_{k}$ for each class $k\in\{1,2,\dots,K\}$, such that $\alpha_{k}=e_{k}+1$. The belief mass [10] $b_{k}$ for each class and the overall uncertainty $u$ are computed as follows:
$b_{k}=\frac{e_{k}}{S};\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ u=\frac{K}{S}$ (3)
where $S=\sum_{i=1}^{K}(e_{i}+1)$. This formulation ensures that the uncertainty $u$ is inversely proportional to the total evidence $S$. When no evidence in support of a prediction is available, the model gives maximum uncertainty ($u=1$).
The learning process involves training a neural network to predict the parameters of the Dirichlet distribution $\boldsymbol{\alpha}$, which are parameterized by the network weights $\Theta$. An evidential network is trained on a classification task by optimizing Bayes risk function that incorporates both prediction accuracy and its uncertainty. This approach is based on empirical findings from [36]:
$\mathcal{L}_{i}(\Theta)$ $=\int\left\|\mathbf{y}_{i}-\mathbf{p}_{i}\right\|_{2}^{2}\frac{1}{B\left(\boldsymbol{\alpha}_{i}\right)}\prod_{i=1}^{K}p_{ij}^{\alpha_{ij}-1}d\mathbf{p}_{i}$ (4)
$=\sum_{j=1}^{K}\mathbb{E}\left[y_{ij}^{2}-2y_{ij}p_{ij}+p_{ij}^{2}\right]$ (5)
$=\sum_{j=1}^{K}\left(y_{ij}^{2}-2y_{ij}\mathbb{E}\left[p_{ij}\right]+\mathbb{E}\left[p_{ij}^{2}\right]\right)$ (6)
The loss function is designed to minimize the prediction error while encouraging the model to produce low-variance distributions. We explore four possible ways of embedding Dirichlet uncertainties into the category-level image retrieval framework, described in the following sections.
2.1 Evidential Classification as a Deep Metric Learning Baseline
Building on the methodology presented in [36], we train DeiT-S [40] and GC ViT Tiny models for the evidential classification task using the CUB-200-2011 dataset. This approach reinterprets the networks’ outputs and introduces a novel loss function, described in Equation 6. By leveraging the CLS token from the Vision Transformers as an image embedding, we perform image retrieval by comparing the embeddings of the query image with those of the database images using cosine similarity. The retrieval process is illustrated in Figure 1.
2.2 $\boldsymbol{\alpha}$-embeddings: Using Dirichlet Distribution Parameters as Neural Codes
In this experiment, instead of using the CLS Token, we employ the final-layer outputs of the network, which now represent the parameters $\boldsymbol{\alpha}$ of the Dirichlet distribution rather than the softmax probability distribution $\mathbf{p}$. These $\boldsymbol{\alpha}$ parameters are utilized directly as image embeddings and are compared using the $L_{2}$ distance metric. This approach aims to evaluate the effectiveness of $\boldsymbol{\alpha}$ parameters as neural representations. The diagram in Figure 1 also applies to this retrieval paradigm.
2.3 Uncertainty-Driven Retrieval Reranking
In this setup, we utilize a contrastively-trained GC ViT model for feature extraction. Subsequently, an evidential GC ViT is employed to derive uncertainty values for each of the top $N$ results. These results are then reranked in ascending order of uncertainty. The process is depicted in Figure 3.
2.4 Distribution-Based Distances
Given that the network outputs now represent the parameters of the Dirichlet distribution, which encode the uncertainty of the network’s predictions, this approach explores the feasibility of directly embedding images into distributions. Instead of comparing vector embeddings, we directly compare these distributional embeddings using the Bhattacharyya distance [33]. Images that are similar are close in distribution space, indicated by a small Bhattacharyya distance between them. The method is illustrated in Figure 2.
Figure 1: Retrieval based on image embeddings. CLS token of the vision transformer (or $\alpha$ vector) is taken as the image embedding, and is used for matching between image pairs using cosine similarity.
Figure 2: Retrieval based on continuous distribution embeddings. Images are mapped to $\alpha$-parameterized Dirichlet distributions, after which matching is performed based on negative Bhattacharyya distance between them, instead of the inner product.
Figure 3: Uncertainty-based re-ranking: First retrieval is performed the regular way, by comparing the image descriptors using the cosine similarity. However, after that, a separate, evidential neural network computes uncertainties for top K results, after which re-ranking based on these uncertainties is performed.
3 Experiements and Results
First we perform experiments which concern the choice of the best-performing backbone architecture for embedding the images for further use in retrieval. Following [12], all models used in this paper are fine-tuned either on a classification task or on the metric-learning task using entropy-regularization. Global Context Vision Transformer has shown to outperform the DeiT-S used in [12], on both the CUB-200-2011 and SOP datasets, by a wide margin, as demonstrated in Table 1, Figure 5, Figure 6.
This led us to adopt the GC ViT as the backbone architecture, and conduct further experiments on the CUB-200-2011 dataset to explore the effectiveness of proposed probabilistic approaches. Results are summarized in Table 2. Evidential classification excelled, surpassing regular classification in terms of the Recall@K metric. The best performing method was Uncertainty-Driven Reranking, while $\boldsymbol{\alpha}$-embeddings and Distribution-embedding methods fell short, yielding results below the baseline.
Table 1: Comparison between the contrastively-trained DeiT-S and GC ViT-T models in category-level image retrieval tasks on the Stanford Online Products and CUB-200-2011 datasets.
Recall@K [%]
Dataset K DeiT-S GC ViT-T
SOP 1 83.28 85.45  [+2.17%]
10 93.08 94.38  [+1.30%]
100 96.86 97.60  [+0.74%]
1000 98.86 99.16  [+0.30%]
CUB 1 62.64 70.80  [+13.02%]
2 73.21 79.54  [+8.64%]
4 81.87 86.85  [+6.08%]
8 88.47 92.40  [+4.44%]
Figure 4: Cosine similiarities of positive and negative pairs produced by two architectures. Greater separation observed with the GC ViT indicates better processing of both local and global relations within images.
Figure 5: Number of queries exceeding a given average precision on the CUB-200-2011 dataset. GC ViT outperforms the DeiT across all average precision thresholds.
Figure 6: Number of queries exceeding a given average precision on the SOP dataset. GC ViT outperforms the DeiT across all average precision thresholds.
Table 2: Evidential retrieval. Recall@K on the CUB-200 2011 dataset, achieved with the GC ViT Tiny model, trained under the various evidential network frameworks. Each row corresponds to a different method of retrieval.
Recall@K [%]
Retrieval type / K 1 2 4 8
Classification (base) 72.89 82.65 88.47 92.94
Evidential class. 80.22 86.38 90.23 93.15
Uncertainty rerank. 80.79 86.09 90.42 93.17
$\boldsymbol{\alpha}$-embeddings 53.87 66.26 77.08 85.04
Distribution emb. 4.00 4.12 5.33 5.98
4 Conclusion
Our research advances the field of content-based image retrieval by introducing and leveraging evidential uncertainty quantification methods. Specifically, we utilize evidential learning to enhance general image retrieval quality, we introduce a general task-agnostic reranking scheme based on uncertainty values, benchmarks novel transformer architectures, and establish evidential classification as a good baseline for metric learning. Our findings suggest that incorporating uncertainty estimates into image retrieval systems can substantially improve their quality, reliability and interpretability. Future research could further explore adversarial robustness, different distribution-based embeddings as well as other probabilistic methods to build upon the foundational steps we established.